# HisTrackMap: Global Vectorized High-Definition Map Construction via History Map Tracking

### v2

@@ 1 Introduction
Figure 1: Comparison of Different HD Map construction Paradigms. (a) Single-Frame, (b) Temporal, (c) HisTrackMap (Ours). The gray square represents detect queries, while the purple square represents temporal propagation queries.
Recent advancements in Perspective View (PV)-to-Bird’s-Eye View (BEV) methods have significantly enhanced vectorized HD map construction such as [10, 14, 13, 7]. These approaches leverage the DETR-based detection paradigm [2] to achieve precise HD map generation. To illustrate the differences across various paradigms, we present a comparison in Fig.1. Specifically, Fig.1 (a) showcases the single-frame architecture, which directly generates vectorized maps by decoding parallel learnable or prior-initialized queries using a decoder to interact with BEV feature. Nevertheless, internal prediction instabilities within the model coupled with uncontrollable environmental factors, such as occlusions or low-light conditions, frequently result in temporal perception inconsistencies, posing substantial challenges for real-world autonomous driving scenarios. Fig.1 (b) describes the temporal architecture, where latent query embedding is used as a stream memory [29, 37], facilitating the propagation of temporal information within a unified latent memory. Furthermore, MapTracker [3] employs a tracking paradigm that further utilizes query propagation and implicit latent memory to associate instance-level map elements across consecutive frames, enhancing temporal consistency. In such a framework, a Motion MLP proposed in [37] is essential for obtaining the next-frame latent memory considering the pose transformation between consecutive frames. Consequently, it employs a transformation loss to guide the Motion MLP, implicitly ensuring the accuracy of the pose transformation.
It is understood that a substantial amount of historical perception map data generated during vehicle navigation could potentially be leveraged as prior knowledge or utilized to reduce computational overhead in overlapping map regions. However, the current approach of implicit encoding through instance queries remains inadequate for precisely recording or preserving past geometric map information. To fully utilize existing perception results, as shown in Fig. 1 (c), we propose HisTrackMap, which explicitly maintains instance-level history maps for corresponding map instances under a tracking paradigm. This approach provides prior information for subsequent navigation perception, ensuring smoother and more continuous HD map construction and enhancing the efficiency and accuracy of temporal information propagation.
@@ 2 Related Work
Figure 3: The architecture pipeline of the HisTrackMap. Initialization at timestamp 0 and tracking at T and T+1 are illustrated. Track queries associate map elements and construct temporal relationships through the history maps under the tracking paradigm.
@@
Evaluation Metrics. MapVR [39] and MGMap [17] evaluate rasterized results using IoU (Intersection over Union) metric. In contrast, MapTR [13, 14] emphasizes point-to-point correspondence and employs Chamfer Distance to calculate average precision. However, these metrics are primarily effective for evaluating instantaneous performance. Since map perception in practice requires evaluation over continuous sequences, these metrics struggle to measure temporal consistency effectively. To address this limitation, MapTracker [3] introduces the consistency-aware metric (C-mAP), which aims to penalize mismatches in perception results based on the alignment of corresponding instances across frames.
@@ 3.1 Overview Architecture
Fig. 3 illustrates the framework of the proposed HisTrackMap, built on a tracking paradigm [3, 38]. The encoder extracts perception features (BEV feature $\mathbf{F}_{bev}$ and PV feature $\mathbf{F}_{pv}$) from surrounding-view images to localize spatial structures. The input detection query $\mathbf{Q}\in\mathbb{R}^{N_{q}\times C}$ is updated through MapDecoder to produce the coordinates $\mathbf{P}\in\mathbb{R}^{N_{q}\times N_{p}\times 2}$, categories $\mathbf{C}\in\mathbb{R}^{N_{q}\times 3}$ (pedestrian, boundry, divider) and scores $\mathbf{S}\in\mathbb{R}^{N_{q}}$. Here, $N_{q}$ denotes the number of detected queries, $N_{p}$ represents the number of points of a map instance, and $C$ is the feature dimension.
@@ 3.2 Instance-Level History Maps
Subsequently, the perception results at timestamp $t$ are used to update the history map. For the $j$-th instance, if it is identified as a newly born instance (i.e., no corresponding tracking index is found, and its confidence score $\mathbf{S}_{j}^{t}$ exceeds the detect threshold $\tau_{det}$), the rasterization method $\operatorname{Raster}(\cdot)$ [39] is employed to transform the vectorized map representation $\mathbf{P}_{j}^{t}$ into a rasterized format. The rasterized result is scaled by $\mathbf{S}_{j}^{t}$ to initialize its history map. If the instance has been tracked (i.e., some tracking index is associated, and $\mathbf{S}_{j}^{t}$ exceeds the tracking threshold $\tau_{track}$), the corresponding historical map $\mathbf{M}_{i}^{t-1}$ is retrieved based on its tracking index $i$. This history map is then temporally decayed using a decay factor $\lambda$ and updated with the newly rasterized result, yielding the updated history map $\mathbf{M}_{i}^{t}$ as:
$\mathbf{M}_{i}^{t}=\begin{cases}\operatorname{Raster}(\mathbf{P}_{j}^{t})\cdot\mathbf{S}_{j}^{t},&\text{if newborn},\\[6.0pt] \max\big{(}\lambda\cdot\mathbf{M}_{i}^{t-1},\operatorname{Raster}(\mathbf{P}_{j}^{t})\cdot\mathbf{S}_{j}^{t}\big{)},&\text{if tracked}.\end{cases}$ (1)
After completing the update process, the history map $\mathbf{M}_{i}^{t}$ is finalized. At timestamp $t$, both the total number of tracked instances $N_{track}^{t}$ and the collection of history maps $\mathcal{M}^{t}$ are updated accordingly. At timestamp t+1, $\mathcal{M}^{t}$ is dynamically aligned with the updated ego position to obtain warped history maps $\mathcal{M}_{w}^{t+1}$, thereby enabling initialization for the subsequent timestamp as
$\mathcal{M}_{w}^{t+1}=\operatorname{Warp}(\mathcal{M}^{t},\mathcal{T}^{t+1}),$ (2)
where $\mathcal{T}^{t+1}$ denotes a standard $4\times 4$ transformation matrix that characterizes the spatial transformation relationship between the coordinate systems of two consecutive frames, $\operatorname{Warp}(\cdot)$ is the alignment function.
Note that, if the prediction confidence of some track query is below the $\tau_{track}$, it indicates that that instance has disappeared or has not been successfully tracked at the current frame. Consequently, it is necessary to remove the corresponding track query and history map $\mathbf{M}_{r}$ from $\mathcal{M}^{t+1}$, and the total number of tracked instances $N_{track}^{t+1}$ will be decremented accordingly as:
$\mathcal{M}^{t+1}=\mathcal{M}_{w}^{t+1}\backslash\{\mathbf{M}_{r}\}.$ (3)
@@ 3.3 Map-Trajectory Prior Fusion
As aforementioned, to address the limitations of implicit temporal propagation in capturing detailed feature transformations during motion, the Map-Trajectory Prior Fusion is designed to integrate historical map trajectory information into track queries, enhancing temporal prior utilization and perceptual consistency. As shown in Fig. 4, the Map-Trajectory Prior Fusion first samples instance features from the history map within the PV and BEV feature spaces. Since the number of instance features varies, padding is applied to align them to a uniform length. These instance-level perception features are then individually integrated into the track queries through cross-attention.
@@
Historical information provides both semantic category and spatial trajectory coordinate data. To enhance the semantic representation of track queries, we define an initial class embedding $\mathbf{CE}_{init}\in\mathbb{R}^{3\times C}$, which encodes previous track categories into a class embedding $\mathbf{CE}\in\mathbb{R}^{N_{track}\times C}$. The class embedding is fused with the track query to provide category priors as
$\mathbf{Q}_{track}=\mathbf{Q}_{track}+\mathbf{CE}.$ (4)
@@
The history map is a fine-grained map with temporal decay. Therefore, we generate the valid pixel mask $\mathcal{M}_{val}$ by filtering pixels in $\mathcal{M}$ that exceed the map threshold $\tau_{map}$. To enhance the spatial awareness of PV features in 3D space, learnable coordinate embeddings $\mathbf{PE}_{coords}\in\mathbb{R}^{h\times w\times C}$ are combined with camera embeddings $\mathbf{PE}_{cam}\in\mathbb{R}^{cams\times C}$ generated via the ego-to-image matrix $\mathbf{E2I}\in\mathbb{R}^{cams\times 4\times 4}$, producing the final PV position embedding $\mathbf{PE}_{pv}\in\mathbb{R}^{cams\times h\times w\times C}$ as
$\mathbf{PE}_{pv}=\mathbf{PE}_{cam}+\mathbf{PE}_{coords},$ (5)
where $h,w$ are the resolutions of the PV features, and $cams$ is the number of cameras.
@@
Then, we project $\mathcal{M}_{val}$ into the perspective view space using the projection function $\operatorname{Proj}(\cdot)$. In this space, the PV features $\mathbf{F}_{pv}$ are combined with the corresponding positional embedding $\mathbf{PE}_{pv}$ to form enhanced feature representations. Subsequently, the $\operatorname{SampledPV}(\cdot)$ function is applied to sample from these enhanced features, resulting in the final sampled PV features $\mathbf{F}_{sampled\_pv}$ as
$\mathbf{F}_{sampled\_pv}=\operatorname{SampledPV}(\operatorname{Proj}(\mathcal{M}_{val},\mathbf{E2I}),\mathbf{F}_{pv}+\mathbf{PE}_{pv}).$ (6)
@@
Similarly, we perform analogous operations in the Bird’s-Eye View space. To ensure that the sampled BEV features $\mathbf{F}_{sampled\_bev}$ incorporate positional information, we introduce a sinusoidal position embedding $\mathbf{PE}_{bev}\in\mathbb{R}^{H\times W\times C}$. Then, we utilize $\mathcal{M}_{val}$ to sample BEV feature $\mathbf{F}_{bev}$ through the $\operatorname{SampledBEV}(\cdot)$ function as
$\mathbf{F}_{sampled\_bev}=\operatorname{SampledBEV}(\mathcal{M}_{val},\mathbf{F}_{bev}+\mathbf{PE}_{bev}).$ (7)
@@
Next, the PV and BEV features corresponding to each instance, including the historical map trajectories, have been obtained. The track queries, along with their associated sampled features $\mathbf{F}_{sampled\_bev}$ and $\mathbf{F}_{sampled\_pv}$, are then utilized within a cross-attention $\operatorname{CA}(\cdot)$ to finalize the initialization of the track queries for the current timestamp:
$\begin{array}[]{l}\mathbf{Q}_{track}=\mathbf{CA}(\mathbf{Q}_{track},\mathbf{F}_{sampled\_pv}),\\[6.0pt] \mathbf{Q}_{track}=\mathbf{CA}(\mathbf{Q}_{track},\mathbf{F}_{sampled\_bev}).\\ \end{array}$ (8)
@@
1:  $\textbf{TP}\leftarrow\text{zeros}(\mathbf{CM}.shape[0])$
2:  $\textbf{FP}\leftarrow\text{zeros}(\mathbf{CM}.shape[0])$
@@ 3.4 Global Geometric HD Mapping Evaluation
Mask2Map [4] [ECCV2024] 24 70.6 71.3 72.9 71.6 55.8 44.5 9.2
MapTracker [3] [ECCV2024] 24 75.3 69.2 71.2 71.9 63.4 47.3 10.9
HisTrackMap (Ours) 24 76.9 72.7 71.9 73.8 64.7 48.5 10.3
@@ 4.1 Experimental Settings
MapTRv2 [14] [IJCV2024] 24 62.9 72.1 67.1 67.4 -- --
HRMapNet [41] [ECCV2024] 30 65.1 71.4 68.6 68.3 -- --
HIMap [43] [CVPR2024] 24 72.4 72.4 73.2 72.7 -- --
MGMapNet [35] [ICLR2025] 24 71.3 76.0 73.1 73.6 -- --
MapTracker [3] [ECCV2024] 24 75.0 78.7 73.1 75.6 63.1 47.4
HisTrackMap (Ours) 24 77.4 78.9 74.1 76.8 66.6 48.2
@@
Table 2: Comparison with SOTA methods on the Argoverse2 validation set. The ’--’ symbol indicates that no results are provided.
@@ 4.4 Ablation and Robustness Studies
The proposed Map-Trajectory Prior Fusion consists of three main components: Class Embedding, PV Prior, and BEV Prior. To evaluate the contributions of each component, we conducted an ablation study on the nuScenes dataset, with the results presented in Tab. 4. The experiment without Map-Trajectory Prior Fusion is used as the baseline. There are several observations from the results. First, introducing Class Embedding improved mAP by 0.7, demonstrating the effectiveness of semantic category priors. Second, incorporating PV Prior and BEV Prior based on Class Embedding further improves mAP from 72.3 to 72.9 and 73.3, respectively. This reveals that it is beneficial to utilize geometric trajectory information to enhance track queries. Third, with the complete Map-Trajectory Prior Fusion, the model achieved the highest 73.8 mAP, highlighting the significance of history map based trajectory priors in effectively leveraging temporal-spatial information to enhance perception performance.
Map-Trajectory Prior Fusion mAP
Class Embed. PV Prior BEV Prior
-- -- -- 71.6
$\checkmark$ -- -- 72.3
$\checkmark$ $\checkmark$ -- 72.9
$\checkmark$ -- $\checkmark$ 73.3
$\checkmark$ $\checkmark$ $\checkmark$ 73.8
Table 4: Ablation study of the Map-Trajectory Prior Fusion.
$\sigma_{t}$ (m)
0 0.1 0.2
$\sigma_{r}$ (rad) 0 73.8 73.6 73.5
0.01 73.5 73.4 73.1
0.02 73.3 73.2 73.0
Table 5: Robustness study with different localization errors.
The update of the history map primarily relies on the vehicle’s motion pose parameters. To evaluate the robustness of HisTrackMap to localization errors, we conducted additional experiments on the nuScenes dataset, as detailed in Tab. 5. Random noise was introduced to the translation and rotation of the extrinsic matrix to disrupt the history map update. The results show that the model achieves 73.0 mAP under noise levels of 0.2 m and 0.02 rad and 73.4 mAP under real-world noise levels (0.1 m, 0.01 rad), maintaining a clear advantage. The experimental results continue to outperform most previous single-frame models, demonstrating the practical effectiveness of HisTrackMap.

### v3

@@ 1 Introduction
Recent advancements in Perspective View (PV)-to-Bird’s-Eye View (BEV) methods have significantly enhanced vectorized HD map construction such as [10, 14, 13, 7]. These approaches leverage the DETR-based detection paradigm [2] to achieve precise HD map generation. To illustrate the differences across various paradigms. Nevertheless, internal prediction instabilities within the model coupled with uncontrollable environmental factors, such as occlusions or low-light conditions, frequently result in temporal perception inconsistencies, posing substantial challenges for real-world autonomous driving scenarios. Latent query embedding is used as a stream memory [29, 37], facilitating the propagation of temporal information within a unified latent memory. Furthermore, MapTracker [3] employs a tracking paradigm that further utilizes query propagation and implicit latent memory to associate instance-level map elements across consecutive frames, enhancing temporal consistency. In such a framework, a Motion MLP proposed in [37] is essential for obtaining the next-frame latent memory considering the pose transformation between consecutive frames. Consequently, it employs a transformation loss to guide the Motion MLP, implicitly ensuring the accuracy of the pose transformation.
It is understood that a substantial amount of historical perception map data generated during vehicle navigation could potentially be leveraged as prior knowledge or utilized to reduce computational overhead in overlapping map regions. However, the current approach of implicit encoding through instance queries remains inadequate for precisely recording or preserving past geometric map information. To fully utilize existing perception results, we propose HisTrackMap, which explicitly maintains instance-level history maps for corresponding map instances under a tracking paradigm. This approach provides prior information for subsequent navigation perception, ensuring smoother and more continuous HD map construction and enhancing the efficiency and accuracy of temporal information propagation.
@@ 3.1 Overview Architecture
The encoder extracts perception features (BEV feature $\mathbf{F}_{bev}$ and PV feature $\mathbf{F}_{pv}$) from surrounding-view images to localize spatial structures. The input detection query $\mathbf{Q}\in\mathbb{R}^{N_{q}\times C}$ is updated through MapDecoder to produce the coordinates $\mathbf{P}\in\mathbb{R}^{N_{q}\times N_{p}\times 2}$, categories $\mathbf{C}\in\mathbb{R}^{N_{q}\times 3}$ (pedestrian, boundry, divider) and scores $\mathbf{S}\in\mathbb{R}^{N_{q}}$. Here, $N_{q}$ denotes the number of detected queries, $N_{p}$ represents the number of points of a map instance, and $C$ is the feature dimension.
@@ 3.2 Instance-Level History Maps
Subsequently, the perception results at timestamp $t$ are used to update the history map. For the $j$-th instance, if it is identified as a newly born instance (i.e., no corresponding tracking index is found, and its confidence score $\mathbf{S}_{j}^{t}$ exceeds the detect threshold $\tau_{det}$), the rasterization method $\operatorname{Raster}(\cdot)$ [38] is employed to transform the vectorized map representation $\mathbf{P}_{j}^{t}$ into a rasterized format. The rasterized result is scaled by $\mathbf{S}_{j}^{t}$ to initialize its history map. If the instance has been tracked (i.e., some tracking index is associated, and $\mathbf{S}_{j}^{t}$ exceeds the tracking threshold $\tau_{track}$), the corresponding historical map $\mathbf{M}_{i}^{t-1}$ is retrieved based on its tracking index $i$. This history map is then temporally decayed using a decay factor $\lambda$ and updated with the newly rasterized result, yielding the updated history map. After completing the update process, the history map $\mathbf{M}_{i}^{t}$ is finalized. At timestamp $t$, both the total number of tracked instances $N_{track}^{t}$ and the collection of history maps $\mathcal{M}^{t}$ are updated accordingly. At the $t+1$ timestamp, if the confidence score exceeds the threshold $\tau_{track}$, the instance will continue to be tracked. At timestamp t+1, $\mathcal{M}^{t}$ is dynamically aligned with the updated ego position to obtain warped history maps $\mathcal{M}_{w}^{t+1}$, thereby enabling initialization for the subsequent timestamp.
Note that, if the prediction confidence of some track query is below the $\tau_{track}$, it indicates that that instance has disappeared or has not been successfully tracked at the current frame. Consequently, it is necessary to remove the corresponding track query and history map $\mathbf{M}_{r}$ from $\mathcal{M}^{t+1}$, and the total number of tracked instances $N_{track}^{t+1}$ will be decremented accordingly.
@@ 3.3 Map-Trajectory Prior Fusion
As aforementioned, to address the limitations of implicit temporal propagation in capturing detailed feature transformations during motion, the Map-Trajectory Prior Fusion is designed to integrate historical map trajectory information into track queries, enhancing temporal prior utilization and perceptual consistency. The Map-Trajectory Prior Fusion first samples instance features from the history map within the PV and BEV feature spaces. Since the number of instance features varies, padding is applied to align them to a uniform length. These instance-level perception features are then individually integrated into the track queries through cross-attention.
@@
Historical information provides both semantic category and spatial trajectory coordinate data. To enhance the semantic representation of track queries, we define an initial class embedding $\mathbf{CE}_{init}\in\mathbb{R}^{3\times C}$, which encodes previous track categories into a class embedding $\mathbf{CE}\in\mathbb{R}^{N_{track}\times C}$.
@@
The history map is a fine-grained map with temporal decay. Therefore, we generate the valid pixel mask $\mathcal{M}_{val}$ by filtering pixels in $\mathcal{M}$ that exceed the map threshold $\tau_{map}$.
@@
Then, we project $\mathcal{M}_{val}$ into the perspective view space using the projection function $\operatorname{Proj}(\cdot)$. In this space, the PV features $\mathbf{F}_{pv}$ are combined with the corresponding positional embedding $\mathbf{PE}_{pv}$ to form enhanced feature representations. Subsequently, the $\operatorname{SampledPV}(\cdot)$ function is applied to sample from these enhanced features, resulting in the final sampled PV features
@@
Similarly, we perform analogous operations in the Bird’s-Eye View space. To ensure that the sampled BEV features $\mathbf{F}_{sampled\_bev}$ incorporate positional information, we introduce a sinusoidal position embedding $\mathbf{PE}_{bev}\in\mathbb{R}^{H\times W\times C}$. Then, we utilize $\mathcal{M}_{val}$ to sample BEV feature $\mathbf{F}_{bev}$ through the $\operatorname{SampledBEV}(\cdot)$ function.
@@
Next, the PV and BEV features corresponding to each instance, including the historical map trajectories, have been obtained. The track queries, along with their associated sampled features $\mathbf{F}_{sampled\_bev}$ and $\mathbf{F}_{sampled\_pv}$, are then utilized within a cross-attention $\operatorname{CA}(\cdot)$ to finalize the initialization of the track queries for the current timestamp.
@@ 4.1 Experimental Settings
MapTRv2 [14] [IJCV2024] 24 62.9 72.1 67.1 67.4 - -
HRMapNet [40] [ECCV2024] 30 65.1 71.4 68.6 68.3 - -
HIMap [42] [CVPR2024] 24 72.4 72.4 73.2 72.7 - -
MGMapNet [35] [ICLR2025] 24 71.3 76.0 73.1 73.6 - -
@@
Table 2: Comparison with SOTA methods on the Argoverse2 validation set. The ’-’ symbol indicates that no results are provided.
@@ 4.4 Ablation and Robustness Studies
The proposed Map-Trajectory Prior Fusion consists of three main components: Class Embedding, PV Prior, and BEV Prior. To evaluate the contributions of each component, we conducted an ablation study on the nuScenes dataset. The experiment without Map-Trajectory Prior Fusion is used as the baseline. There are several observations from the results. First, introducing Class Embedding improved mAP by 0.7, demonstrating the effectiveness of semantic category priors. Second, incorporating PV Prior and BEV Prior based on Class Embedding further improves mAP from 72.3 to 72.9 and 73.3, respectively. This reveals that it is beneficial to utilize geometric trajectory information to enhance track queries. Third, with the complete Map-Trajectory Prior Fusion, the model achieved the highest 73.8 mAP, highlighting the significance of history map based trajectory priors in effectively leveraging temporal-spatial information to enhance perception performance.
The update of the history map primarily relies on the vehicle’s motion pose parameters. To evaluate the robustness of HisTrackMap to localization errors, we conducted additional experiments on the nuScenes dataset. Random noise was introduced to the translation and rotation of the extrinsic matrix to disrupt the history map update. The results show that the model achieves 73.0 mAP under noise levels of 0.2 m and 0.02 rad and 73.4 mAP under real-world noise levels (0.1 m, 0.01 rad), maintaining a clear advantage. The experimental results continue to outperform most previous single-frame models, demonstrating the practical effectiveness of HisTrackMap.